OpenStudy (anonymous):

Put each equation into slope intercept form y=3x-4 and y=-2x+11 Aren't these both already in slope intercept form?

OpenStudy (anonymous):

yes its already in slope intercept form :)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Ok thanks so how do I put this equation into slope intercept form? 4x-y=1 and 5x+2=24 ?

OpenStudy (anonymous):

What about 5x+2=24 ?

OpenStudy (anonymous):

5x+2=24 5x=24-2 x= 22/5 no slope or slope is 0

OpenStudy (anonymous):

4x-y=1 y=4x-1 slope is 4

OpenStudy (anonymous):

Sorry I forgot the y it should be 5x+2y=24

OpenStudy (anonymous):

so then would it be y=2.5x+12 ?

OpenStudy (anonymous):

So I don't divide the 24 by 2 also ? which would be y=-2.5x+12 ?

OpenStudy (anonymous):

sorry I caculator wrong 5x+2y=24 2y = 24-5x y= (24-5x)/2 y = -2.5x+12 yes, but slope is negative

OpenStudy (anonymous):

Ok that helps thank you!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

Can you work a couple more with me so I make sure I have this right ? y=-3x+6 ( is already in slope intercept form 2y=10x-36

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

So would it be y=5x-18 ?

OpenStudy (anonymous):

y=-3x+6 ( is already in slope intercept form yes

OpenStudy (anonymous):

y=5x-18 correct

OpenStudy (anonymous):

Ok one more y=-x-9 ( already in slope intercept form) 3y=5x+5

OpenStudy (anonymous):

y=-x-9 ( already in slope intercept form) yes

OpenStudy (anonymous):

y=1.6x+1.6 ?

OpenStudy (anonymous):

3y=5x+5 y=5x/3 +5/3

OpenStudy (anonymous):

So then do you divide the 5/3 to get 1.6 or leave it in the fraction form?

OpenStudy (anonymous):

if decimal leave in fraction\[y=\frac{5}{3}x+\frac{5}{3}\]